OpenStudy (anonymous):

The length of a rectangle is twice the width. The area is 162yd^2. Find the length and the width. The length is_____yd The width is_____yd

OpenStudy (lgbasallote):

l = 2w area = l x w = 162 so \[2w \times w = 162\] \[2w^2 = 162\] divide by 2 \[w^2 = 81\] can you do it from there?

OpenStudy (anonymous):

No im still a little confused

OpenStudy (lgbasallote):

get the square root of both sides

OpenStudy (lgbasallote):

what is the square root of 81?

OpenStudy (anonymous):

9

OpenStudy (lgbasallote):

therefore width is 9 it is said that length is two times the width so what is 2 x 9

OpenStudy (anonymous):

18?

OpenStudy (lgbasallote):

width = 9 length = 18 dont forget units...what is the unit?

OpenStudy (anonymous):

Is it 2? I'm confused

OpenStudy (lgbasallote):

units is meters, kilometers, those things

OpenStudy (lgbasallote):

what is the unit of measurement for this one?

OpenStudy (anonymous):

yards right

OpenStudy (lgbasallote):

yep yards

OpenStudy (anonymous):

Thanks, I have to go we are having sever weather...